# DPoser: Diffusion Model as Robust 3D Human Pose Prior

### 09 Dec 2023 v1

@@ 6 Conclusion
We introduce DPoser, a robust and vetsatile human pose prior framework grounded in diffusion models. Engineered for flexibility, DPoser shines across multiple pose-related tasks, enhanced by our innovative truncated timestep scheduling for test-time optimization. Comprehensive experiments substantiate DPoser’s superior performance over existing state-of-the-art pose priors, establishing it as an effective solution for a wide range of applications.
Limitation and future work. While our framework benefits from variational diffusion sampling [18], it also shares its limitations, such as the mode-seeking behavior. Future research could look into enhancing solution diversity via techniques like particle-based variational inference [30, 31]. Furthermore, within the broader context of inverse problems, a plethora of existing methods [22, 23, 25, 24] could be adapted to leverage our diffusion-based prior. Exploring these methods holds great potential for future progress.
Supplementary Material:
DPoser: Diffusion Model as Robust 3D Human Pose Prior
@@ A Parameterization of Score-based Diffusion Models
The resulting form of Eq. (16) aligns precisely with the noise prediction form of diffusion models [10] (refer to Eq. (4)). This implies that by training $\epsilon_{\theta}(\mathbf{x}_{t};t)$ in a diffusion model context, we simultaneously get a handle on the score function, approximated as $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$.
@@ B View DPoser as Score Distillation Sampling
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
1step 57.28 84.79 39.77 36.15 38.21 19.87
5steps 57.61 85.14 41.30 37.58 40.22 21.21
10steps 57.66 85.07 41.28 37.39 40.69 21.34
@@
Interestingly, the gradient of DPoser in Eq. (10) coincides with Score Distillation Sampling (SDS) [27, 28], which can be interpreted as aiming to minimize the following KL divergence:
@@
where $p_{t}^{\mathtt{SDE}}\left(\mathbf{x}_{t};\theta\right)$ denote the marginal distribution whose score function is estimated by $\epsilon_{\theta}(\mathbf{x}_{t};t)$. For the specific case where $t\to 0$, this term encourages the Dirac distribution $\delta(\mathbf{x}_{0})$ (i.e., the optimized variable) to gravitate toward the learned data distribution $p_{0}^{\mathtt{SDE}}\left(\mathbf{x}_{0};\theta\right)$, while the Gaussian perturbation like Eq. (17) softens the constraint. Building on this understanding, we can borrow advanced techniques from SDS [27, 28]—a rapidly evolving area ripe for methodological innovations [31, 32, 29]. To extend this, we experiment with a multi-step denoising strategy adapted from HiFA [29], substituting our original one-step denoising process. This alternative, however, yields suboptimal results across all evaluation metrics, as demonstrated in Tab. S-1. A plausible explanation could be that our proposed truncated timestep scheduling effectively manages low noise levels (i.e., small $t$), thus negating the need for more denoising steps.
@@ C Experimental Details
This section elaborates on the specifics of our pose completion and motion denoising experiments, and also highlights DPoser’s proficiency in pose interpolation tasks.
@@ C.1 Pose completion
For partial observations $\mathbf{y}$, the measurement operator $\mathcal{A}$ is modeled as a mask matrix $M\in\mathbb{R}^{d\times n}$. Based on our optimization framework (i.e., Algorithm 1), we define the task-specific loss, $L_{\text{comp}}$, as follows:
@@
Here, $\mathbf{x}_{0}$ denotes the whole body pose $\theta$ we try to recover, where the unseen parts are initialized as random noise. In ourthe following ablated studies, if not specified, the evaluation is performed using 10 hypotheses on the AMASS [52] test set with left leg occlusion.
C.2 Motion denoising
@@
In complement to the comparative analysis presented in Tab. 4, we extend our evaluation to include scenarios with varying noise levels. This extended examination, detailed in Tab. S-2, showcases DPoser’s exceptional performance against state-of-the-art (SOTA) pose priors, especially under conditions of high noise, manifesting DPoser’s resilience to noise.
C.3 Pose interpolation
(a) Pose reconstruction.
(b) Interpolation between two poses.
Figure S-1: Visual results of pose reconstruction and interpolation.
For the forward SDE (Eq. (1)), there exists a probability flow ODE with the same marginal probability densities corresponding to it, expressed as: [8]
$\mathrm{d}\mathbf{x}=[\mu(t)\mathbf{x}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)]\mathrm{d}t.$ (21)
By replacing $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ with our approximation $-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$, this ODE can be treated as a neural ODE [5], which is fully invertible. By integrating through it, we can encode any data point $\mathbf{x}_{0}$ into its latent representation $\bar{\mathbf{x}}_{1}$ and decode it through its reverse-time counterpart.
For interpolation, we start with two poses, $\mathbf{x}_{0}^{0}$ and $\mathbf{x}_{0}^{1}$, and encode them into latent representations $\bar{\mathbf{x}}_{1}^{0}$ and $\bar{\mathbf{x}}_{1}^{1}$. Utilizing spherical linear interpolation (Slerp), we obtain interpolated latents $\bar{\mathbf{x}}_{1}^{t}=\text{Slerp}(\bar{\mathbf{x}}_{1}^{0},\bar{\mathbf{x}}_{1}^{1};t)$ for $t\in[0,1]$, which are then decoded to produce novel poses. The Slerp function is defined as:
$\text{Slerp}(A,B;t)=\frac{\sin((1-t)\omega)}{\sin\omega}A+\frac{\sin(t\omega)}{\sin\omega}B,$ (22)
where $\omega=\arccos(A\cdot B/\|A\|\|B\|)$.
To validate the informativeness of our encoding, we employ MPJPE and MPVPE metrics for reconstruction evaluation. Compared to VPoser [35], a VAE-based prior, DPoser exhibits superior performance with 4.79mm and 6.58mm errors, against 9.85mm and 12.16mm, respectively. Qualitative results further confirm DPoser’s strong reconstruction and smooth interpolation capabilities, as shown in Fig. S-1.
D Ablated DPoser Training
@@
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
w/o norm 58.62 83.75 60.70 55.47 44.82 24.04
min-max 59.55 85.63 58.18 53.06 42.70 21.29
z-score 57.66 85.47 52.86 48.27 38.57 20.24
Table S-3: Evaluation of DPoser’s performance under different normalization methods, specifically for the axis-angle rotation representation.
@@
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
axis-angle 57.28 84.79 52.14 47.60 38.21 19.87
6D rotations 58.02 85.93 56.28 51.54 38.44 20.12
Table S-4: Comparative performance of rotation representations under z-score normalization across multiple tasks and metrics.
Figure S-2: Evolution of MPJPE during DPoser training, evaluated on pose completion. The evaluation is conducted on the AMASS [52] test set, utilizing 10 hypotheses and simulating occlusion scenarios affecting both legs.
Building on the insights shared in Sec. 4.1 in our main text, this part dissects the impact of different rotation representations and normalization techniques on DPoser’s performance. Initially, we examine axis-angle representation, comparing various normalization strategies: min-max scaling, z-score normalization, and no normalization. Our findings, summarized in Tab. S-3, indicate that z-score normalization is generally the most effective. Subsequently, using this optimal normalization, we explore 6D rotations [59] as an alternative. As evidenced by Tab. S-4, axis-angle representation offers superior performance. This preference can be attributed to the effective modeling capabilities of diffusion models, along with the inherent advantages of axis-angle in capturing bounded joint rotations.
Inspired by HuMoR [39], we experiment with integrating the SMPL body model [33] as a regularization term during DPoser’s training. Alongside the prediction of additive noise, as outlined in Eq (4), we employ a 10-step DDIM sampler [11] to recover a “clean” version of the pose, denoted as $\tilde{\mathbf{x}}_{0}$, from the diffused $\mathbf{x}_{t}$. The regularization loss aims to minimize the discrepancy between the original and recovered poses under the SMPL body model $M$:
$L_{\mathrm{reg}}=$ $||M_{J}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{J}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}$
$+||M_{V}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{V}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}.$
Here, $\beta_{0}$ represents the mean shape parameters in SMPL. To account for denoising errors, we scale the regularization loss by $\mathrm{log}(1+\frac{\alpha_{t}}{\sigma_{t}})$, thereby increasing the weight for samples with smaller $t$ values (less noise). Fig. S-2 visualizes the impact of this regularization on MPJPE during the training, specifically for pose completion tasks with occlusion of both legs. We observe that weighted regularization offers slight performance gains in the early training process, while the absence of weighting introduces instability and deterioration in results. Despite these insights, the computational cost of incorporating the SMPL model—especially for our large batch size of 1280—makes the training approximately 8 times slower. Therefore, we opted not to include this regularization in our main experiments.
E Extended DPoser Optimization
@@
MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
ScoreSDE [8] 52.86 48.27 80.68 74.74 82.38 65.81 27.97 17.27
DPS [19] 41.99 38.08 66.16 61.28 64.75 51.75 25.68 15.90
MCG [20] 51.97 47.46 80.07 74.14 82.22 65.58 27.93 17.27
DPoser (ours) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
Table S-5: Comparative evaluation of diffusion-based solvers for pose completion.
In addressing pose-centric tasks as inverse problems, we propose a versatile optimization framework, which employs variational diffusion sampling as its foundational approach [18]. Our exploration extends to an array of diffusion-based methodologies for solving these complex inverse problems. Among the techniques considered are ScoreSDE [8], MCG [20], and DPS [19]. These methods augment standard generative processes with observational data, either by employing gradient-based guidance or back-projection techniques. We compare these methods with our DPoser for pose completion tasks. Our findings, captured in Tab. S-5, reveal that DPoser outperforms the competitors under most occlusion conditions. Consequently, DPoser emerges not merely as a universally applicable solution to pose-related tasks, but also as an exceptionally efficient one.
@@
F More Qualitative Results
We show more qualitative results for pose generation (Fig. S-3), pose completion (Fig. S-4), human mesh recovery (Fig. S-5), and motion denoising (Fig. S-6).

### 23 Mar 2024 v2

@@ 5 Conclusion
We introduce DPoser, to our best knowledge, the first unconditional diffusion-based pose prior, tailored for an expansive array of pose-related tasks. Engineered for flexibility, DPoser can be implemented as a straightforward L2-loss regularizer and enhanced by our innovative truncated timestep scheduling for test-time optimization. Comprehensive experiments substantiate DPoser’s superior performance over existing state-of-the-art pose priors.
Limitation and future work. While our framework benefits from variational diffusion sampling [33], it also shares its limitations, such as the mode-seeking behavior. Future research could look into enhancing solution diversity via techniques like particle-based variational inference [30, 53]. Furthermore, within the broader context of inverse problems we have framed, a plethora of existing methods [45, 2, 5, 35] could be adapted to leverage our diffusion-based prior. Exploring these methods holds great potential for future progress.
Ethical Considerations. For a discussion on the potential negative impacts of our work, please refer to the Appendix.
Appendix for DPoser: Diffusion Model as
Robust 3D Human Pose Prior
In this appendix, we first briefly recap the parameterization of diffusion models and their connection to score functions in Sec. A, followed by the perspective of Score Distillation Sampling (SDS) to understand our DPoser regularization in Sec. B. We detail the experimental setup and nuances in Sec. C and dissect various training aspects of DPoser in Sec. D. The exploration of extended optimization techniques is discussed in Sec. E, and considerations for truncated timestep scheduling in image domains are presented in Sec. F. Additional qualitative results are showcased in Sec. G. Lastly, potential negative impacts such as biases in data and ethical concerns in application are considered in Sec. H.
@@ A Parameterization of Score-based Diffusion Models
The resulting form of Eq. (16) aligns precisely with the noise prediction form of diffusion models [16] (refer to Eq. (4) in the main text). This implies that by training $\epsilon_{\theta}(\mathbf{x}_{t};t)$ in a diffusion model context, we simultaneously get a handle on the score function, approximated as $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$.
@@ B View DPoser as Score Distillation Sampling
PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
1 step 56.05 35.37/74.01/26.47 38.21 19.87
5 steps 56.16 36.59/80.82/31.22 40.22 21.21
10 steps 56.18 36.78/82.59/32.32 40.69 21.34
@@
Interestingly, the gradient of DPoser (Eq. (10) in the main text) coincides with Score Distillation Sampling (SDS) [38, 52], which can be interpreted as aiming to minimize the following KL divergence:
@@
where $p_{t}^{\mathtt{SDE}}\left(\mathbf{x}_{t};\theta\right)$ denote the marginal distribution whose score function is estimated by $\epsilon_{\theta}(\mathbf{x}_{t};t)$. For the specific case where $t\to 0$, this term encourages the Dirac distribution $\delta(\mathbf{x}_{0})$ (i.e., the optimized variable) to gravitate toward the learned data distribution $p_{0}^{\mathtt{SDE}}\left(\mathbf{x}_{0};\theta\right)$, while the Gaussian perturbation like Eq. (17) softens the constraint. Building on this understanding, we can borrow advanced techniques from SDS [38, 52]—a rapidly evolving area ripe for methodological innovations [53, 54, 58]. To extend this, we experiment with a multi-step denoising strategy adapted from HiFA [58], substituting our original one-step denoising process. This alternative, however, yields suboptimal results across all evaluation metrics, as demonstrated in Tab. S-1. A plausible explanation could be that our proposed truncated timestep scheduling effectively manages low noise levels (i.e., small $t$), thus negating the need for more denoising steps. In addition, iterative denoising in each optimization step may cause error accumulations, leading to inaccurate gradients.
@@ C Experimental Details
This section elaborates on the specifics of our pose completion and motion denoising experiments.
@@ C.1 Pose Completion
For partial observations $\mathbf{y}$, the measurement operator $\mathcal{A}$ is modeled as a mask matrix $M\in\mathbb{R}^{d\times n}$. Based on our optimization framework (Algorithm 1 in the main text), we define the task-specific loss, $L_{\text{comp}}$, as follows:
@@
Here, $\mathbf{x}_{0}$ denotes the complete body pose $\theta$ we try to recover, where the unseen parts are initialized as random noise. In the following ablated studies, if not specified, the evaluation is performed using 10 hypotheses on the AMASS [32] dataset with left leg occlusion.
C.2 Motion Denoising (Noisy Input)
@@
In complement to the comparative analysis presented in Table 4 of our main text, we extend our evaluation to include scenarios with varying noise levels. This extended examination, detailed in Tab. S-2, showcases DPoser’s exceptional performance against state-of-the-art (SOTA) pose priors, especially under conditions of high noise, manifesting DPoser’s resilience to noise.
C.3 Motion Denoising (Partial Input)
This task focuses on reconstructing clean poses, $\theta^{t}$, from partially observed joint positions, $J_{\text{obs}}^{t}$, across $N$ frames, employing a known mask matrix to identify visible joints. The optimization objective mirrors that of motion denoising (Sec. C.2), but incorporates a mask in Eq. (19) to specifically target visible parts, ensuring that only these segments guide the recovery process.
We conducted experiments on the AMASS dataset [32] to assess our model’s performance on this task with two types of occlusions: legs and left arm. The quantitative results of these experiments are detailed in Tab. S-3, and the accompanying visualizations are provided in Sec. G.
In leg occlusion scenarios, the AMASS dataset primarily showcases straight poses, offering minimal diversity. This scenario permits decent outcomes without incorporating a pose prior, since the optimization’s starting point closely aligns with these prevalent poses. However, VPoser’s mean-centered characteristic hinders its ability to faithfully replicate the visible areas. On the other hand, Pose-NDF falls short in enhancing the occluded parts. DPoser accurately handles visible parts and guides occluded ones for more realistic poses. For left arm occlusions, which involve more varied movements, DPoser markedly surpasses other methods, underlining its adaptability and precision in handling diverse motion patterns.
Methods Occlusion MPJPE MPVPE
Vis. Occ. All. All.
No prior Legs 0.26 14.72 5.52 5.45
VPoser Legs 1.75 14.29 6.31 7.38
PoseNDF Legs 0.25 15.71 5.87 5.64
DPoser (ours) Legs 0.28 12.24 4.63 3.65
No prior Left Arm 0.26 24.87 4.74 9.91
VPoser Left Arm 1.21 13.23 3.40 7.68
PoseNDF Left Arm 0.25 17.70 3.42 7.86
DPoser (ours) Left Arm 0.27 7.80 1.64 3.81
Table S-3: Comparative analysis of methods for motion denoising with different occlusions (Legs and Left Arm) on the AMASS dataset. Errors (in cm) are evaluated in terms of MPJPE across visible (Vis.), occluded (Occ.), and all joints, along with MPVPE for all vertices.
D Ablated DPoser’s Training
@@
PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
w/o norm 57.88 45.37/102.28/41.08 44.82 24.04
min-max 59.17 47.41/107.00/43.42 42.70 21.29
z-score 56.49 34.37/72.47/26.32 38.57 20.24
Table S-4: Evaluation of DPoser’s performance under different normalization methods, specifically for the axis-angle rotation representation.
@@
PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
axis-angle 56.05 34.76/72.41/26.09 38.21 19.87
6D rotations 57.54 40.89/81.43/27.31 38.44 20.12
Table S-5: Comparative performance of rotation representations under z-score normalization across multiple tasks and metrics.
This section dissects the impact of different rotation representations and normalization techniques on DPoser’s performance. Initially, we examine axis-angle representation, comparing various normalization strategies: min-max scaling, z-score normalization, and no normalization. Our findings, summarized in Tab. S-4, indicate that z-score normalization is generally the most effective. Subsequently, using this optimal normalization, we explore 6D rotations [57] as an alternative. As evidenced by Tab. S-5, axis-angle representation offers superior performance. This preference can be attributed to the effective modeling capabilities of diffusion models, along with the inherent advantages of axis-angle in capturing bounded joint rotations for regression tasks like human mesh recovery.
Inspired by HuMoR [40], we experiment with integrating the SMPL body model [31] as a regularization term during training. Alongside the prediction of additive noise, as outlined in Equation (4) in the main text, we employ a 10-step DDIM sampler [44] to recover a “clean” version of the pose, denoted as $\tilde{\mathbf{x}}_{0}$, from the diffused $\mathbf{x}_{t}$. The regularization loss aims to minimize the discrepancy between the original and recovered poses under the SMPL body model $M$:
$L_{\mathrm{reg}}=||M_{J}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{J}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}+||M_{V}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{V}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}.$ (21)
Here, $\beta_{0}$ represents the mean shape parameters in SMPL. To account for denoising errors, we scale the regularization loss by $\mathrm{log}(1+\frac{\alpha_{t}}{\sigma_{t}})$, thereby increasing the weight for samples with smaller $t$ values (less noise).
Fig. S-1 visualizes the impact of this regularization on MPJPE during the training, specifically for pose completion tasks with occlusion of both legs.
Figure S-1: MPJPE evolution in DPoser training for pose completion, assessed on AMASS [32] with 10 hypotheses under legs occlusion scenarios.
We observe that weighted regularization offers slight performance gains in the early training process, while the absence of weighting introduces instability and deterioration in results. Despite these insights, the computational cost of incorporating the SMPL model—especially for our large batch size of 1280—makes the training approximately 8 times slower. Therefore, we opted not to include this regularization in our main experiments.
E Extended DPoser’s Optimization
@@
ScoreSDE [48] 48.73/106.32/41.30 74.68/128.32/37.27 66.89/127.86/48.15 16.69/34.54/12.21
DPS [6] 40.51/104.32/54.57 64.26/113.46/33.71 60.63/119.85/42.78 15.10/33.90/13.27
MCG [7] 49.04/106.37/41.07 74.90/128.53/37.40 66.17/127.72/48.15 16.69/34.66/12.23
DPoser(ours) 35.37/74.01/26.47 59.25/96.77/24.55 51.27/81.76/20.04 13.95/28.57/9.85
Table S-6: Comparative evaluation of diffusion-based solvers for pose completion on the AMASS dataset [32] (hypotheses number $S=10$).
In addressing pose-centric tasks as inverse problems, we propose a versatile optimization framework, which employs variational diffusion sampling as its foundational approach [33]. Our exploration extends to an array of diffusion-based methodologies for solving these complex inverse problems. Among the techniques considered are ScoreSDE [48], MCG [7], and DPS [6]. These methods augment standard generative processes with observational data, either by employing gradient-based guidance or back-projection techniques. We compare these methods with our DPoser for pose completion tasks. Our findings, captured in Tab. S-6, reveal that DPoser outperforms the competitors under most occlusion conditions. Consequently, DPoser emerges not merely as a universally applicable solution to pose-related tasks, but also as an exceptionally efficient one.
@@
F Truncated Timestep Scheduling on Images
Figure S-2: Image inpainting using standard (a) and truncated (b) timestep scheduling. The process evolution is shown over iterations with the middle row depicting the log-magnitude spectrum and the bottom row the phase spectrum.
Exploring truncated timestep scheduling for image-based tasks, we find its suitability for human poses doesn’t translate well to images. Initial timesteps are critical in image domains for generating foundational perceptual content.
In our study, we employed a 256x256 unconditional diffusion model [11] trained on ImageNet [10] with variational diffusion sampling [33] for image inpainting. Comparing standard (timesteps 990 to 0) and truncated scheduling (timesteps 495 to 0), both with 100 steps, the experiments confirmed that truncation compromises image quality (Fig. S-2). The standard approach preserved perceptual content, while truncation produced disjointed patches, misaligned with the original image context.
These results affirm that truncated timestep scheduling excels in pose data where key information emerges in later stages but falls short in image tasks where early timesteps are essential. This scheduling is thus bespoke to the characteristics of human pose estimation and is unsuitable for image processes that rely on the full diffusion timeline for content fidelity.
G More Qualitative Results
We show more qualitative results for pose generation (Fig. S-3), pose completion (Fig. S-4), human mesh recovery (Fig. S-5) and motion denoising (Fig. S-6, Fig. S-7).
H Potential Negative Impacts
  • Bias and Fairness Concerns: Human pose prior learning models may inadvertently encode biases present in the training data, leading to biased predictions or discriminatory outcomes. This can perpetuate existing societal biases and inequalities, particularly if the training data is not representative or balanced across diverse demographics.
  • Ethical Considerations: The use of human pose prior learning models in applications such as surveillance, security, or healthcare raises ethical concerns regarding individual privacy, autonomy, and consent. There are debates about the appropriate use of such technologies and the potential for unintended consequences or misuse.
  • Dependency on Data Quality: Human pose prior learning models heavily rely on the quality and diversity of the training data. Poorly annotated or biased datasets can negatively impact the performance and reliability of these models, leading to inaccurate or unreliable predictions.